# On Some Quantum Correction to the Coulomb Potential in Generalized Uncertainty Principle Approach

### 07 Jan 2024 v1

@@ IV Solutions of the Coulomb(-4) model with a GUP
where the new function $\varphi_{n}^{G}(x)$ is a polynomial function and the parameters $a>0\,,b\,,c\,,d,\,f>0$ are still unknown, but must be such that the original function $\psi_{n}^{G}(x)$ is square integrable. Substituting (33) into (3) and solving the Riccati equation that results for $g(x)$, it can be seen that the previous parameters are determined by
@@
$\left\{P_{4}(x)\frac{d^{2}}{dx^{2}}+Q_{4}(x)\frac{d}{dx}+W_{3}(x)\right\}\varphi_{n}^{GUP}(x)=0,$ (35)
@@
the parameters $a\,,b\,,c\,,d,\,f$ and $\gamma_{i}\,$, $i=1,2,3,4$ are given respectively by (IV) and (II). To solve (35), we use the general Bethe ansatz method introduced in Appendix A in Ref. AgZh2013 . In this way, we look for polynomial solutions for $\varphi_{n}^{G}(x)$ of the form
@@
where $x_{i}$ are distinct roots to be determined. Applying then the Eqs. (A.6)-(A.10) from AgZh2013 , the ‘general solutions’ of (35) are given by this set of equations:
@@
Note that the term in square brackets is positive since $\epsilon_{n}<0$ times (21), again confirming that $\alpha_{1}$ must be negative for (40) to be satisfied. Through simple manipulations, (40) can be rewritten in closed form:
$\epsilon_{n}^{G}(\epsilon_{n};\alpha_{1},\beta)=-\frac{\alpha_{1}^{2}}{4}\left(\frac{2\beta\,\epsilon_{n}-1}{\alpha_{1}\,\sqrt{\beta}+(n+2)}\right)^{2}+\beta\,\epsilon_{n}^{2}\,.$ (41)
Let us keep in mind that the denominator of the fraction, $\alpha_{1}\,\sqrt{\beta}+(n+2)$, is always positive because $f>0$, as established in (IV), and remember that $\epsilon_{n}$ is given by (21). On the other hand, the equations (38b)-(38d) produce severe restrictions on the potential parameters, explicitly:
@@
$\sum_{j=1,\,j\neq i}^{n}\frac{1}{x_{i}-x_{j}}-\frac{x_{i}^{4}\sqrt{\beta\,\epsilon_{n}^{2}-\epsilon_{n}^{G}}-\sqrt{\beta}\left(\alpha_{4}+\alpha_{1}\,x_{i}^{3}+\alpha_{2}\,x_{i}^{2}+\alpha_{3}\,x_{i}\right)-2x_{i}^{3}}{x_{i}^{4}}=0,\quad i=1,2,\dots,n.$ (43)
In summary, for a given $n$, the ‘general solutions’ of the Schrödinger equation modified with a GUP are given by the equations (33) together with (37) and (41)-(43) As in the case of the ordinary model (without GUP), below we will look for explicit solutions for the ground state and the first excited state.
@@
For $n=0$, from (41), the ground-state energy, $\epsilon_{0}^{G}$, is determined in closed form
@@
where the ground-state energy of the ordinary case, $\epsilon_{0}$, was already obtained explicitly in (22). The explicit form of the associated wave function, from (33) together with (IV) and (37), is given by
@@
where $C_{0}$ is the normalization constant and, due to (IV), the potential parameters satisfy the constraints
@@
It is interesting to note that the Bethe ansatz (39) equations play no role in the ground state solution. In this simple case, substituting $\alpha_{2}$ and $\alpha_{3}$ into $\alpha_{4}$ in (IV), we find that $-\frac{2}{3\sqrt{\beta}}<\alpha_{1}<0$ so that $\alpha_{4}>0$ according to the assumption made in (5).
@@
provided $\alpha_{1}<0$. Recall again that $\epsilon_{1}$ was already obtained in (25). In the present case it is assumed that the differential equation (35) has solutions in the form (37), that is, $\varphi_{1}^{GUP}(x)=x-x_{1}\,$, and consequently, from (33)–(IV), the wave function is explicitly given by
@@
The unknown parameter $x_{1}$ appearing in (48) and (IV) is determined from the Bethe ansatz equation (39), i.e.
@@ V Conclusions
Based on recent papers that propose a correction of the Coulomb potential by including quadratic, cubic and inverse quartic terms, we consider this modified potential in the framework of the modified Schrödinger equation with a generalized uncertainty principle. The resulting equation appears as a generalization of the Heun equation, that is, the biconfluent Heun equation that has the form $y^{\prime\prime}(x)+\left(A_{0}+\frac{A_{-1}}{x}+\frac{A_{-2}}{x^{2}}+\frac{A_{-3}}{x^{3}}+\frac{A_{-4}}{x^{4}}\right)y(x)=0$ Ronveaux . As far as we know, the general form we considered, could not be solved by other analytical techniques, including the common Lie-algebraic approach, which requires the form $P_{4}(x)\Phi^{\prime\prime}(x)+Q_{3}(x)\Phi^{\prime}(x)+W_{2}(x)\Phi(x)=0$ Turbiner ; Turbiner88 . Therefore, we try a Bethe-Ansatz to the equation and after some calculations we find the ground state and the first excited state. It is worth mentioning that despite its merits, this approach becomes longer for higher states as the parameters involved and their determination become more complicated, as some constraints are imposed on the parameters in order to provide the solutions. However, the truth is that this last limitation exists in most approaches, including supersymmetric quantum mechanics, power series solutions, integral transforms, etc. We are currently working on the generalization of these solutions to few-body systems, to problems in more dimensions and to other physical fields.
@@ Acknowledgments
L.M.N. and S.Z. were supported by Spanish MCIN with funding from European Union Next Generation EU (PRTRC17.I1) and Consejeria de Educacion from JCyL through QCAYLE project, as well as grants PID2020-113406GB-I00 MTM funded by AEI/10.13039/501100011033, and RED2022-134301-T. The work of M.B. was supported by the Czech Science Foundation within the project 22-18739S.

### 21 Mar 2024 v2

@@ IV Solutions of the Coulomb–4 model with a GUP
where the new function $\varphi_{n}^{G}(x)$ must be a polynomial and the parameters $a>0\,,b\,,c\,,d>0,\,f>0$ are still unknown, but must be such that the original function $\psi_{n}^{G}(x)$ is square integrable. Substituting (36) into (3) and solving the Riccati equation that results for $g(x)$, it can be seen that the previous parameters are determined by
@@
$\left\{P_{4}(x)\frac{d^{2}}{dx^{2}}+Q_{4}(x)\frac{d}{dx}+W_{3}(x)\right\}\varphi_{n}^{G}(x)=0,$ (38)
@@
the parameters $a\,,b\,,c\,,d,\,f$ and $\gamma_{i}\,$, $i=1,2,3,4$ are given respectively by (IV) and (II). To solve (38) we use the general Bethe ansatz method, introduced in Appendix A of Ref. AgZh2013 . In this way, we look for polynomial solutions for $\varphi_{n}^{G}(x)$ of the form
@@
where $x_{i}$ are distinct roots to be determined. Now applying Eqs. (A.6)–(A.10) of AgZh2013 , the ‘general solutions’ of (38) are given by this set of equations:
@@
Note that the term in square brackets is positive since, according to (21), $\epsilon_{n}<0$, again confirming that $\alpha_{1}$ must be negative for (43) to hold. Through simple manipulations, (43) can be rewritten in closed form like this:
$\epsilon_{n}^{G}(\epsilon_{n};\alpha_{1},\beta)=-\frac{\alpha_{1}^{2}}{4}\left(\frac{2\beta\,\epsilon_{n}-1}{\alpha_{1}\,\sqrt{\beta}+n+2}\right)^{2}+\beta\,\epsilon_{n}^{2}\,.$ (44)
Let us keep in mind that the denominator of the fraction, $\alpha_{1}\,\sqrt{\beta}+n+2$, is always positive because $f>0$, as established in (IV), and remember that $\epsilon_{n}$ is given by (21). On the other hand, the equations (41b)–(41d) produce severe restrictions on the potential parameters, explicitly:
@@
$\sum_{j=1,\,j\neq i}^{n}\frac{1}{x_{i}-x_{j}}-\frac{x_{i}^{4}\sqrt{\beta\,\epsilon_{n}^{2}-\epsilon_{n}^{G}}-\sqrt{\beta}\left(\alpha_{4}^{2}+\alpha_{1}\,x_{i}^{3}+\alpha_{2}\,x_{i}^{2}+\alpha_{3}\,x_{i}\right)-2x_{i}^{3}}{x_{i}^{4}}=0,\quad i=1,2,\dots,n.$ (46)
In summary, for a given $n$, the ‘general solutions’ of the Schrödinger equation modified with a GUP are given by the equations (36) together with (40) and (44)–(46). As in the ordinary case (without a GUP), we will now look for explicit solutions for the ground state and the first excited state.
@@
For $n=0$, from (44) the energy of ground state, $\epsilon_{0}^{G}$, is determined in closed form:
@@
where the ground state energy of the ordinary case, $\epsilon_{0}$, was already obtained in (22). From (36) together with (IV) and (40), the explicit form of the associated wave function is given by
@@
where $C_{0}$ is the normalization constant and, due to (IV), the parameters of the potential satisfy the constraints
@@
Note that the Bethe ansatz equations (42) play no role in the ground state solution given the fact that $\varphi_{0}^{G}(x)\equiv 1$, cf. (40). It can also be seen that the above restrictions imply that $\alpha_{2}$ and $\alpha_{3}$ are positive in the case of GUP, taking into account that $\epsilon_{0}<0$. As a consequence, from the equation (49a), we find that $-\frac{2}{3\sqrt{\beta}}<\alpha_{1}<0$.
@@
provided $\alpha_{1}<0$. Recall again that $\epsilon_{1}$ was already obtained in (25). In the present case it is assumed that the differential equation (38) has solutions in the form of (40), that is, $\varphi_{1}^{G}(x)=x-x_{1}\,$, and consequently, from (36)–(IV), the wave function is explicitly given by
@@
The unknown parameter $x_{1}$, which appears in (51) and (IV), is determined from the Bethe ansatz equation (42), which turns out to be
@@ V Conclusions
Based on recent papers proposing a correction of the Coulomb potential by including inverse quadratic, cubic and quartic terms, we consider this modified potential in the framework of the modified Schrödinger equation with a generalized uncertainty principle. The resulting equation appears as a generalization of the Heun equation, more precisely, the biconfluent Heun equation that has the form $y^{\prime\prime}(x)+\left(A_{0}+\frac{A_{-1}}{x}+\frac{A_{-2}}{x^{2}}+\frac{A_{-3}}{x^{3}}+\frac{A_{-4}}{x^{4}}\right)y(x)=0$ Ronveaux . To the best of our knowledge, the general form we consider could not be solved by other analytical techniques, including the common Lie algebraic approach, which requires the form $P_{4}(x)\Phi^{\prime\prime}(x)+Q_{3}(x)\Phi^{\prime}(x)+W_{2}(x)\Phi(x)=0$ Turbiner ; Turbiner88 . Therefore, we attempted a Bethe analysis of the equation and after some calculations we found the general solution to the modified problem due to the GUP and then both the ground state and the first excited state are explicitly determined. It is worth mentioning that, despite its merits, this approach is very burdensome for higher states, since the determination of the parameters involved becomes more complicated, due to the restrictions that must be imposed on the parameters in order to provide solutions. However, the truth is that this last limitation exists in most approaches, including supersymmetric quantum mechanics, power series solutions, integral transforms, etc.
Let us now review our solutions to realize that we have already given the solutions to some important physical problems. The first application to consider is the simple but rich toy model of the mini-universe proposed by Roshan Roshan to study the expansion (or contraction) of a homogeneous and isotropic universe. Equation (27) from Roshan’s paper, which is the governing quantum equation, can be converted to that of our problem by a transformation $\phi(x)=R(x)\psi(x)$. Now, having calculated the solutions of the ordinary cases of the so-called generalized Coulomb–4 potential, all the results there, including the mass and the Bohr radius, can now be derived simply with a quite realistic potential, which includes both post-Newtonian relativistic and quantum corrections.
A second application of our work is to the so-called Vaz quantum gravitational model. In a very recent paper Corda , Corda reviews the firewall paradox and Vaz’s quantum gravitational model Corda . The Vaz quantum mechanical model was originally formulated to study the spherical collapse of inhomogeneous dust and is AdS of dimension $d=n+2$. The author considers a two-body Hamiltonian and, having separated the problem into a center-of-mass framework, derives a Schrödinger-type equation with an inverse term plus a gravitational potential, which is also Coulomb-type. The potential considered there is the ordinary gravitational potential, the energy of which determines the mass and energy spectra of the black hole and the density of the Vaz shell. Based on the results obtained here for the ordinary case of the one-dimensional Schrödinger equation, one can simply revise that model with the generalized gravitational potential with a simple parameter change in the calculations.
A third similarity to the structure analyzed in the present work arises with a semi-relativistic approximation of a quantum field theoretical two-body Bethe-Salpeter equation, which is often called the spinless Salpeter equation, and finds applications in a variety of physical fields Lucha ; Adamo .
A fourth category of problems that is related to our work is that of wave equations in curved space or wave equations in the space-time of cosmic strings, such as, for example, the three-dimensional Dirac equation in a rather general curved spacetime Jahangiri ; Medeiros . It should be noted that in the similar cases mentioned, there is not as much difficulty as here in treating the associated ordinary energy levels as a direct parameter in the calculation.
In conclusion, we think that the present work provides a basis for addressing some important problems in other related fields and that it could be useful for further studies. In particular, a generalization of the GUP with this interaction to a three-dimensional case and a two-body formulation will be a suitable framework for future theoretical investigations or probably for designing experimental setups. The latter could be particularly attractive in relation to Planck-scale physics and quantum optics Pikovski , since a direct and realistic quantum term is now available, which needs further consideration as to its exact value.
@@ Acknowledgments
The research of L.M.N. and S.Z. was supported by the European Union.-Next Generation UE/MICIU/Plan de Recuperacion, Transformacion y Resiliencia/Junta de Castilla y Leon, RED2022-134301-T financed by MICIU/AEI/10.13039/501100011033, and PID2020-113406GB-I00 financed by MICIU/AEI/10.13039/501100011033. The work of M.B. was supported by the Czech Science Foundation within the project 22-18739S.